# A New Crossover Algorithm for LP Inspired by the Spiral Dynamic of PDHG

### v1

@@ 1 Introduction
Motivated by applications of large-scale LP, there has been a recent trend towards developing first-order methods (FOMs) for LP, such as PDLP (Applegate et al., 2021; Lu and Yang, 2023a; Lu et al., 2023), SCS (O’Donoghue et al., 2016; O’Donoghue, 2021), and ABIP (Lin et al., 2021; Deng et al., 2024). The key advantage of FOMs is their low iteration cost, primarily involving matrix-vector multiplications, which avoids the need for solving linear equations as required in the simplex methods and IPMs. This advantage makes them particularly well-suited for GPU implementation. Among these, PDLP stands out as the most promising, with its GPU implementation demonstrating numerical performance on par with state-of-the-art LP solvers on standard benchmark sets and demonstrating superior performance on large-scale instances (Lu and Yang, 2023a; Lu et al., 2023). PDLP is based on the primal-dual hybrid gradient (PDHG) method, enhanced by various numerical improvements (Applegate et al., 2021). Variants of PDLP have been implemented in both commercial and open-source optimization solvers, such as COPT (Ge et al., 2024), Xpress (Xpress, 2014), Google OR-Tools (Perron and Furnon, 2024), HiGHS (Huangfu and Hall, 2018), and Nvidia cuOpt (Fender, 2024).
@@ 1.1 Related Work
PDHG was studied in Chambolle and Pock (2011) for image processing applications. Applegate et al. (2021) developed a restarted PDHG algorithm for solving linear programs (LPs), which, together with a few practical enhancements, leads to the PDLP solver. Compared with other popular FOM solvers such as SCS and ABIP, there is no linear system to solve in PDLP, making it a more suitable choice for large-scale LPs. cuPDLP, the GPU implementation of PDLP, demonstrated strong numerical performance on par with commercial LP solvers, and has attracted attention from both the academic and solver industry (Lu and Yang, 2023a; Lu et al., 2023). Recently, additional numerical enhancements have been proposed to speed up the algorithm further. Xiong and Freund (2024) pointed out that the central-path-based scaling can markedly improve the convergence rate of PDHG. Lu and Yang (2024a) embedded Halpern iteration (Halpern, 1967) in PDLP, achieving accelerated theoretical guarantees and improved computational results.
The practical success of PDLP motivates a stream of theoretical analysis. Lu and Yang (2022) showed PDHG has a linear convergence rate when solving LP. Applegate et al. (2023) revealed that an accelerated linear convergence rate can be achieved with restarts, and such an accelerated linear rate matches the lower bound. Applegate et al. (2024) illustrated that for infeasible/unbounded LP, PDHG iterates diverge towards the direction of infeasibility/unboundedness certificate; thus, one can detect the infeasibility/unboundedness of LP using PDHG without additional effort. Lu and Yang (2024b) discovered that PDHG has a two-stage behavior when solving LP. In the first stage, PDHG uses finite iterations to verify the optimal active variables; in the second stage, PDHG solves a homogeneous linear inequality system with a significantly improved linear rate. This explains the slow behaviors of PDHG when the LP is near-degeneracy. Xiong and Freund (2024) introduced level-set geometry that characterizes the difficulty of PDHG for LP, which shows that PDHG converges faster for LP with a regular non-flat level-set around the optimal solution set.
@@
With the boom of various IPMs and MIP applications, crossover gradually plays a significant role in the LP solvers. The key task becomes refining IPM solutions to optimal vertices. Mehrotra (1991) added a controlled random cost perturbation to eliminate dual degeneracy at the sacrifice of losing only a little optimality. Especially for nondegenerate LPs, an indicator (Tapia and Zhang, 1991) can be calculated from the interior solution to identify the basis. Bixby and Saltzman (1994) implemented a modified simplex method, which accepts an interior feasible point as a super-basic solution. Then similar to a normal simplex method, Bixby’s algorithm uses the primal (or dual) ratio test to push primal (or dual) variables to bound if possible or pivots it into (or out of) the bound. Andersen and Ye (1996) used the strictly complementary property of IPM solutions to define a perturbed LP problem. Andersen’s method obtains a basis and then recovers feasibility by pivoting and taking a weighted average with the given IPM solution. Ge et al. (2021) also proposed a perturbation crossover method that combines variable fixing and random cost perturbation. They proved that as long as the perturbation is sufficiently small, their crossover approach will not bring objective loss.
@@ 2.1.1 PDHG for LP
Output: The $\epsilon$-optimal solution $(\mathbf{x},\mathbf{y})$.
@@ 1 $k\leftarrow 0$ ;
3       $\mathbf{x}^{(k+1)}=\operatorname{Proj}_{\mathbb{R}^{n}_{+}}\left(\mathbf{x}^{(k)}-\eta(\mathbf{c}-\mathbf{A}^{\top}\mathbf{y}^{(k)})\right)$ ;
4       $\mathbf{y}^{(k+1)}=\mathbf{y}^{(k)}+\eta(\mathbf{b}-\mathbf{A}(2\mathbf{x}^{(k+1)}-\mathbf{x}^{(k)}))$ ;
@@ 2.1.2 Infimal Displacement Vector
If $\eta\in(0,\frac{1}{\|\mathbf{A}\|_{2}})$, $\mathbf{T}$ is firmly non-expansive with respect to the norm $\|\cdot\|_{\mathbf{M}}$ where the positive definite matrix $\mathbf{M}=\begin{bmatrix}\mathbf{I}_{n}&\eta\mathbf{A}^{\top}\\ \eta\mathbf{A}&\mathbf{I}_{m}\end{bmatrix}$. Its IDV is defined as the minimum perturbation we should subtract from $\mathbf{T}$ to ensure it has a fixed point:
@@
Let $\mathbf{T}$ be a non-expansive operator and $\mathbf{z}$ is the initial point. Then the infimal displacement vector of the operator $\mathbf{T}$ satisfies
@@ 2.2 PDHG Spiral Behavior
In this part, we formally present the PDHG spiral behavior when solving LPs. At a high level, PDHG identifies different bases from phase to phase. During each phase, the iterates follow a spiral ray, where the ray direction and the spiral center are derived later.
@@
Notice that PDHG has already found the optimal PDHG basis in Phase 2, but changes the basis two more times due to the obstruction from the lower bound of $\mathbf{x}_{2}$. Actually, Phase 2 and Phase 4 share the same spiral center, and the directions of their spiral rays are both zero vectors.
Another interesting fact is that no basis change event will happen if we remove the bounds $\mathbf{x}_{1},\mathbf{x}_{2}\geq 0$. Phase 1 will continue and be fully observed as a combination of rotation and forward movement, as displayed in Figure 2(b). Moreover, the rotation converges rapidly; thus, PDHG may diverge almost straightly along the IDV, which is an instance of the spiral ray.
Note that the basis change events split the PDHG trajectory into multiple similar spiral phases. Our next goal is to analyze the spiral behavior within each phase, when the PDHG basis remains unchanged.
@@ 2.2.3 Basis Change Event
It is worth highlighting that differing from the optimality improvement in one phase, the basis change event is usually triggered by the combined action of rotation and forward movement, instead of by one side alone. For example, though the spiral center may lie outside the bounds as shown in Figure 3(a), the spiral ray is likely to intersect the bounds, leading to a basis change event similar to the second case in Figure 3(b).
@@ 3 Crossover inspired by PDHG
Proof.
The KKT condition of (1) is
$\mathbf{A}\mathbf{x}=\mathbf{b},\ \mathbf{x}\geq\mathbf{0}$ (15)
$\mathbf{c}-\mathbf{A}^{\top}\mathbf{y}\geq\mathbf{0}$
$\mathbf{c}^{\top}\mathbf{x}-\mathbf{b}^{\top}\mathbf{y}=0.$
@@
Another important phenomenon, which is the key to our crossover method, is that when PDHG starts from a primal (or dual) feasible solution within one phase, it also becomes the start of the primal (or dual) PDHG spiral ray:
@@
Proof.
To prove $\mathbf{x}_{\mathbf{v}}=\mathbf{x}^{(0)}$, we have $(\mathbf{x}_{\mathbf{v}})_{N}=\mathbf{x}^{(0)}_{N}=\mathbf{0}$ and
$=\mathbf{x}^{(0)}_{B}.$
The second equation utilizes the property of Moore–Penrose inverse that $(\mathbf{A}_{B}^{\top}\mathbf{A}_{B})^{\dagger}\mathbf{A}_{B}^{\top}=\mathbf{A}_{B}^{\dagger}=\mathbf{A}_{B}^{\top}(\mathbf{A}_{B}\mathbf{A}_{B}^{\top})^{\dagger}$. Analogously, we also have $\mathbf{y}_{\mathbf{v}}=\mathbf{y}^{(0)}$. ∎
@@ 3.1 Crossover Framework
Many crossover frameworks have been designed (Megiddo, 1991; Bixby and Saltzman, 1994; Andersen and Ye, 1996), most combining simplex methods and IPMs. Following Megiddo’s idea, our crossover mainly includes three parts.
@@
Output: The optimal vertex solution $(\mathbf{x},\mathbf{y})$ and the basis $(B,N)$.
@@
$B=\{i:\mathbf{x}_{i}>0\}$ ;
  // initialize basis set
Solve the auxiliary LP (16) with cost perturbation ;
  // solve auxiliary LP
@@
3       Generate a primal direction $(\bm{\delta}_{\mathbf{x}})_{B}\neq\mathbf{0}$ in the kernel of $\mathbf{A}_{B}$ by solving (17) and calculating the primal spiral ray in (7) with cost perturbation ;
@@
                   $(\bm{\delta}_{\mathbf{x}})_{B}=-(\bm{\delta}_{\mathbf{x}})_{B}$ ;
@@
$D=\{i:\mathbf{c}_{i}-\mathbf{A}_{i}^{\top}\mathbf{y}=0\},N=\{i:\mathbf{c}_{i}-\mathbf{A}_{i}^{\top}\mathbf{y}>0\}$ ;
@@
Solve the auxiliary LP (18) with penalty ;
  // solve auxiliary LP
@@
14       Generate a dual direction $\bm{\delta}_{\mathbf{y}}\neq\mathbf{0}$ in the kernel of $\mathbf{A}_{D}^{\top}$ by solving (19) and calculating the dual spiral ray in (7) with right-hand-side perturbation ;
@@
                   $\bm{\delta}_{\mathbf{y}}=-\bm{\delta}_{\mathbf{y}}$ ;
@@
  // construct the final basis set
@@
The fundamental difference between our crossover algorithm and Megiddo’s scheme is that we use directions $(\bm{\delta}_{\mathbf{x}})_{B}$ and $\bm{\delta}_{\mathbf{y}}$ inspired by PDHG rather than the pivot direction in simplex methods to move the variables. Another difference from Meggido’s framework is that we check the linear independence after the dual push is completed to separate the two processes more independently, while Meggido checks the linear independence immediately once the dual push identifies newly activated dual constraints. In terms of theoretical guarantees, following the same analysis of Megiddo (1991), we can show that this PDHG-inspired crossover will succeed with probability (w.p.) 1.
@@
Our primal push blends two approaches: solving one auxiliary LP and a few sequential OLS problems that identify the ray of PDHG iterates.
After recognizing the basic set $B$, we build an auxiliary LP with random cost perturbation $\tilde{\mathbf{c}}_{B}$,
$\min$ $\tilde{\mathbf{c}}_{B}^{\top}\mathbf{x}_{B}$ (16)
$\operatornamewithlimits{s.t.}$ $\mathbf{A}_{B}\mathbf{x}_{B}=\mathbf{b}$
$\mathbf{x}_{B}\geq\mathbf{0},$
and solve (16) with PDLP. This auxiliary LP has multiple optimal solutions w.p. 0 (see Lemma 1 in Ge et al. (2021)), so with proper perturbation to avoid the unbounded case, we have the unique optimal solution $\mathbf{x}_{B}$ to (16). Then $(\mathbf{x}_{B},\mathbf{0}_{E})$ can be the primal part of an optimal vertex of (1).
Unfortunately, due to the numerical residuals in practice, we may not correctly fix all the variables on bounds at once. To reduce the negative impact of the issue, the first thing to consider is to control the amplitude of disturbance, see Mehrotra (1991); Ge et al. (2021) for more specific discussions. Besides, a verification scheme needs to be designed.
Therefore, after solving (16), we solve a series of OLS problems (17) with various randomly generated costs $\tilde{\mathbf{c}}_{B}$ to calculate the PDHG primal ray directions (7).
@@
From Proposition 10, since the current primal solution $\mathbf{x}_{B}$ is still feasible for (16) with cost perturbation, it is located at the primal spiral center of the rotation in the current phase of PDHG. We can then start from $\mathbf{x}_{B}$ and move along $(\bm{\delta}_{\mathbf{x}})_{B}$ until reaching new boundaries, mirroring how the PDHG solves the auxiliary LPs. At the end of the primal push, no more primal variable can be pushed to its bound when columns of $\mathbf{A}_{B}$ are linearly independent. OLS can be solved by the direct QR factorization, LSMR (Fong and Saunders, 2011), or the conjugate gradient (CG) method.
The reason why we only solve one auxiliary LP is that the auxiliary LP (16) is in most cases enough to obtain a primal solution close to a vertex and that the OLS has lower complexity to reduce the support and verify the termination of primal push. Directly following the primal spiral ray avoids unnecessary oscillation. However, when the original LP is too large, or the resulting OLS is too hard to solve, using multiple auxiliary LPs can be beneficial for the primal push.
@@
In the dual push, we still combine the auxiliary LP and sequential OLS.
For the auxiliary LP, we sum up the non-negative dual slacks as a penalty to activate more dual constraints.
$\min$ $\mathbf{1}^{\top}(\mathbf{c}_{N}-\mathbf{A}_{N}^{\top}\mathbf{y})$ (18)
$\mathbf{A}_{N}^{\top}\mathbf{y}\leq\mathbf{c}_{N}.$
Then similarly, we compute a series of OLS (19) with perturbed right-hand side $\tilde{\mathbf{b}}$ to calculate the dual ray directions of PDHG (7).

### v2

@@ 1 Introduction
Motivated by applications of large-scale LP, there has been a recent trend towards developing first-order methods (FOMs) for LP, such as PDLP (Applegate et al., 2021; Lu and Yang, 2023a; Lu et al., 2023), SCS (O’Donoghue et al., 2016; O’Donoghue, 2021), and ABIP (Lin et al., 2021; Deng et al., 2024). The key advantage of FOMs is their low iteration cost, primarily involving matrix-vector multiplications, which avoids the need for solving linear equations as required in the simplex methods and IPMs. This advantage makes them particularly well-suited for GPU implementation. Among these, PDLP stands out as the most promising, with its GPU implementation demonstrating numerical performance on par with state-of-the-art LP solvers on standard benchmark sets and demonstrating superior performance on large-scale instances (Lu and Yang, 2023a; Lu et al., 2023). PDLP is based on the primal-dual hybrid gradient (PDHG) method, enhanced by various numerical improvements (Applegate et al., 2021). Variants of PDLP have been implemented in both commercial and open-source optimization solvers, such as COPT (Ge et al., 2025a), Xpress (Xpress, 2014), Google OR-Tools (Perron and Furnon, 2024), HiGHS (Huangfu and Hall, 2018), and NVIDIA cuOpt (Blin, 2024).
@@ 1.1 Related Work
PDHG was originally proposed in Zhu and Chan (2008); Esser et al. (2010), and further studied in Chambolle and Pock (2011) for image processing applications. Applegate et al. (2021) developed a restarted PDHG algorithm for solving linear programs (LPs), which, together with a few practical enhancements, leads to the PDLP solver. Compared with other popular FOM solvers such as SCS and ABIP, there is no linear system to solve in PDLP, making it a more suitable choice for large-scale LPs. cuPDLP, the GPU implementation of PDLP, demonstrated strong numerical performance on par with commercial LP solvers, and has attracted attention from both the academic and solver industry (Lu and Yang, 2023a; Lu et al., 2023). Recently, additional numerical enhancements have been proposed to speed up the algorithm further. Xiong and Freund (2024) pointed out that the central-path-based scaling can markedly improve the convergence rate of PDHG. Lu and Yang (2024b) embedded Halpern iteration (Halpern, 1967) in PDLP, achieving accelerated theoretical guarantees and improved computational results.
The practical success of PDLP motivates a stream of theoretical analysis. Lu and Yang (2022) showed PDHG has a linear convergence rate when solving LP. Applegate et al. (2023) revealed that an accelerated linear convergence rate can be achieved with restarts, and such an accelerated linear rate matches the lower bound. Applegate et al. (2024) illustrated that for infeasible/unbounded LP, PDHG iterates diverge towards the direction of infeasibility/unboundedness certificate; thus, one can detect the infeasibility/unboundedness of LP using PDHG without additional effort. Lu and Yang (2024a) discovered that PDHG has a two-stage behavior when solving LP. In the first stage, PDHG uses finite iterations to verify the optimal active variables; in the second stage, PDHG solves a homogeneous linear inequality system with a significantly improved linear rate. This explains the slow behaviors of PDHG when the LP is near-degeneracy. Xiong and Freund (2024) introduced level-set geometry that characterizes the difficulty of PDHG for LP, which shows that PDHG converges faster for LP with a regular non-flat level-set around the optimal solution set.
@@
With the boom of various IPMs and MIP applications, crossover gradually plays a significant role in the LP solvers. The key task becomes refining IPM solutions to optimal vertices. Mehrotra (1991) added a controlled random cost perturbation to eliminate dual degeneracy at the sacrifice of losing only a little optimality. Especially for non-degenerate LPs, an indicator (Tapia and Zhang, 1991) can be calculated from the interior solution to identify the basis. Bixby and Saltzman (1994) implemented a modified simplex method, which accepts an interior feasible point as a super-basic solution. Then similar to a normal simplex method, Bixby’s algorithm uses the primal (or dual) ratio test to push primal (or dual) variables to bound if possible or pivots it into (or out of) the bound. Andersen and Ye (1996) used the strictly complementary property of IPM solutions to define a perturbed LP problem. Andersen’s method obtains a basis and then recovers feasibility by pivoting and taking a weighted average with the given IPM solution. Ge et al. (2025b) also proposed a perturbation crossover method that combines variable fixing and random cost perturbation. They proved that as long as the perturbation is sufficiently small, their crossover approach will not bring objective loss.
@@ 2.1.1 PDHG for LP
Output: The $\epsilon$-optimal solution $(\mathbf{x}^{(k)},\mathbf{y}^{(k)})$.
@@ 1 $k\leftarrow 0$ ;
3 $\mathbf{x}^{(k+1)}\leftarrow\operatorname{Proj}_{\mathbb{R}^{n}_{+}}\left(\mathbf{x}^{(k)}-\eta(\mathbf{c}-\mathbf{A}^{\top}\mathbf{y}^{(k)})\right)$ ;
4 $\mathbf{y}^{(k+1)}\leftarrow\mathbf{y}^{(k)}+\eta(\mathbf{b}-\mathbf{A}(2\mathbf{x}^{(k+1)}-\mathbf{x}^{(k)}))$ ;
@@ 2.1.2 Infimal Displacement Vector
Its IDV is defined as the minimum perturbation we should subtract from $\mathbf{T}$ to ensure it has a fixed point:
@@
If $\eta\in(0,\frac{1}{\|\mathbf{A}\|_{2}})$, $\mathbf{T}$ is firmly non-expansive with respect to the norm $\|\cdot\|_{\mathbf{M}}$ (Applegate et al., 2024), where the positive definite matrix $\mathbf{M}=\begin{bmatrix}\mathbf{I}_{n}&\eta\mathbf{A}^{\top}\\ \eta\mathbf{A}&\mathbf{I}_{m}\end{bmatrix}$. For firmly non-expansive operators, one can recover the IDV using the difference of iterates and/or scaled iterates:
@@
Let $\mathbf{T}$ be a non-expansive operator and $\mathbf{z}$ be the initial point. Then the infimal displacement vector of the operator $\mathbf{T}$ satisfies
@@ 2.2 PDHG Spiral Behavior
In this part, we formally present the PDHG spiral behavior when solving LPs. At a high level, PDHG identifies different bases from phase to phase. During each phase, the iterates follow a spiral ray, where the ray direction and the spiral center have closed-form formulas.
@@
Note that PDHG has already found the optimal PDHG basis in Phase 2, but changes the basis two more times due to the obstruction from the lower bound of $\mathbf{x}_{2}$. Actually, Phase 2 and Phase 4 share the same spiral center, and the directions of their spiral rays are both zero vectors.
Another interesting fact is that no basis change event will happen if we remove the bounds $\mathbf{x}_{1},\mathbf{x}_{2}\geq 0$. Phase 1 will continue and be fully observed as a combination of rotation and forward movement, as displayed in Figure 2(b). Moreover, the rotation converges rapidly; thus, PDHG may diverge almost straight along the IDV, which is an instance of the spiral ray.
The basis change events split the PDHG trajectory into multiple similar spiral phases. Our next goal is to analyze the spiral behavior within each phase, when the PDHG basis remains unchanged.
@@ 2.2.1 PDHG Spiral Behavior within One Phase
To be more specific, within one phase, PDHG classifies $\mathbf{x}$ into $\mathbf{x}_{B}$ and $\mathbf{x}_{N}$, where $\mathbf{x}_{B}$ are movable and $\mathbf{x}_{N}$ are temporarily treated as if fixed at their bounds. $\mathbf{P}_{B}$ is the linear operator that causes rotation. $(\mathbf{x}_{\mathbf{v}})_{B}$ is part of the spiral center corresponding to the movable part of $\mathbf{x}$. $(\mathbf{v}_{\mathbf{x}})_{B}$ is part of the ray direction corresponding to the movable part of $\mathbf{x}$.
@@ 2.2.3 Basis Change Event
It is worth highlighting that, differing from the optimality improvement in one phase, the basis change event is usually triggered by the combined action of rotation and forward movement, instead of by one side alone. For example, though the spiral center may lie outside the bounds as shown in Figure 3(a), the spiral ray is likely to intersect the bounds, leading to a basis change event similar to the second case in Figure 3(b).
@@ 3 Crossover Inspired by PDHG
The proof is given in Appendix C.
@@
Another important phenomenon, which is the key to our crossover method, is that when PDHG starts from a primal (or dual) feasible solution within one phase, it also becomes the start of the primal (or dual) PDHG spiral ray.
@@
The proof is provided in Appendix D.
During each crossover push, Proposition 10 inspires us to guide those unfixed variables along the PDHG ray direction.
@@ 3.1 Crossover Framework
Many crossover frameworks have been designed (Megiddo, 1991; Bixby and Saltzman, 1994; Andersen and Ye, 1996), most combining simplex methods and IPMs. Following Megiddo’s idea, our crossover mainly includes three parts:
@@
[…]
@@
$B\leftarrow\{i:\mathbf{x}_{i}>0\}$ ;
  // initialize basic set
@@
3   Generate a primal direction $(\bm{\delta}_{\mathbf{x}})_{B}\neq\mathbf{0}$ in the kernel of $\mathbf{A}_{B}$ by solving (15) and calculating the primal spiral ray in (7) with cost perturbation and $\eta=1$ ;
@@
       $(\bm{\delta}_{\mathbf{x}})_{B}\leftarrow-(\bm{\delta}_{\mathbf{x}})_{B}$ ;
@@
$D\leftarrow\{i:\mathbf{c}_{i}-\mathbf{A}_{i}^{\top}\mathbf{y}=0\}$, $N\leftarrow\{i:\mathbf{c}_{i}-\mathbf{A}_{i}^{\top}\mathbf{y}>0\}$ ;
@@
14   Generate a dual direction $\bm{\delta}_{\mathbf{y}}\neq\mathbf{0}$ in the kernel of $\mathbf{A}_{D}^{\top}$ by solving (16) and calculating the dual spiral ray in (7) with right-hand-side perturbation and $\eta=1$ ;
@@
       $\bm{\delta}_{\mathbf{y}}\leftarrow-\bm{\delta}_{\mathbf{y}}$ ;
@@
  // construct the optimal basic set
@@
The fundamental difference between our crossover algorithm and Megiddo’s scheme is that we use directions $(\bm{\delta}_{\mathbf{x}})_{B}$ and $\bm{\delta}_{\mathbf{y}}$ inspired by PDHG rather than the pivot direction in simplex methods to move the variables. Another difference from Megiddo’s framework is that we check the linear independence after the dual push is completed to separate the two processes more independently, while Megiddo checks the linear independence immediately once the dual push identifies newly activated dual constraints. In terms of theoretical guarantees, following the same analysis of Megiddo (1991), we can show that this PDHG-inspired crossover will succeed with probability (w.p.) 1.
@@
After recognizing the basic set $B$, we solve a series of OLS subproblems (15) with various randomly generated costs $\tilde{\mathbf{c}}_{B}$ to calculate the PDHG primal ray directions (7).
@@
From Proposition 10, since the current primal solution $\mathbf{x}_{B}$ is still feasible for (1) with cost perturbation, it is located at the primal spiral center of the rotation in the current phase of PDHG. We can then start from $\mathbf{x}_{B}$ and move along $(\bm{\delta}_{\mathbf{x}})_{B}$ until reaching new boundaries, mirroring how the PDHG solves the cost-perturbed LPs. At the end of the primal push, no more primal variable can be pushed to its bound when columns of $\mathbf{A}_{B}$ are linearly independent. OLS can be solved by the direct QR factorization, LSMR (Fong and Saunders, 2011), or the conjugate gradient (CG) method.
@@
In the dual push, similarly, we compute a series of OLS subproblems (16) with perturbed right-hand side $\tilde{\mathbf{b}}$ to calculate the dual ray directions of PDHG (7).
@@
Complexity results.
Our crossover will succeed w.p. 1 without too many OLS subproblems, as stated in Theorem 11. Moreover, the number of OLS subproblems is strongly polynomial w.p. 1.
Theorem 11.
Consider the proposed crossover framework (Algorithm 2). With probability 1, the algorithm terminates and outputs a primal-dual optimal vertex solution with at most $n$ primal and/or dual push steps. Consequently, it requires solving at most $n$ OLS subproblems.
The proof is provided in Appendix E.